Problem 4-22 Break-Even and Target Profit Analysis; Margin of Safety; CM Ratio [LO5-1, LO5-3, LO5-5, LO5-6, LO5-7]

Menlo Company distributes a single product. The company’s sales and expenses for last month follow:

Total    Per Unit
  Sales $ 604,000 $ 40     
  Variable expenses 422,800 28     
  Contribution margin 181,200 $ 12     
  Fixed expenses 146,400
  Net operating income $   34,800

Required:

1. What is the monthly break-even point in unit sales and in dollar sales?

2. Without resorting to computations, what is the total contribution margin at the break-even point?

3-a. How many units would have to be sold each month to earn a target profit of $58,800? Use the formula method.

3-b. Verify your answer by preparing a contribution format income statement at the target sales level.

4. Refer to part 3 and now assume that the tax rate is 30%. How many units would need to be sold each month to an after-tax target profit of $58,800?  (Round the final answer to the nearest whole number.)

5. Refer to the original data. Compute the company's margin of safety in both dollar and percentage terms. (Round your percentage answer to 2 decimal places (i.e .1234 should be entered as 12.34).)

6. What is the company’s CM ratio? If monthly sales increase by $75,000 and there is no change in fixed expenses, by how much would you expect monthly net operating income to increase?

Next Visit question map

Question 4 of 8 Total
Answer #1

1) Break even unit = 146400/12 = 12200 Units

Break even Sales = 12200*40 = $488000

2) Total Contribution margin = 146400

3a) Required unit = (146400+58800)/12 = 17100 Units

3b) Contribution margin income statement

Sale (17100*40) 684000
Variable cost (17100*28) 478800
Contribution margin 205200
Fixed cost 146400
Operating income 58800

4) Income before tax = (58800*100/70) = 84000

Required unit = (146400+84000)/12 = 19200 Units

5) Margin of safety ($) = 604000-488000 = 116000

Margin of safety (%) = 116000/604000 = 19.21%

6) CM ratio = 12/40 = 30%

Net operating income increase by (75000*30%) = $22500